# A Matlab-based Toolbox for Automatic EMT Modeling and Small-Signal Stability Analysis of Modern Power Systems

### v1

@@ Abstract
Index Terms:
Small-signal Analysis, Electromagnetic Transient Model, Hybrid AC/DC System, Impedance Analysis, Passivity Analysis
@@ I-A Motivation
Modern power systems are experiencing a deep transformation as a result of the massive integration of Inverter-Based Resourcess. Power electronics are being used massively in industrial and domestic loads, renewable generation units (mainly solar photovoltaics, onshore and offshore wind), Battery Energy Storage Systems (BESS), Flexible AC transmission systems (FACTS), and High-Voltage-Direct-Current (HVDC) transmission systems. Power electronics-based systems have different characteristics compared to Synchronous Generators, introducing dynamic phenomena on a much faster timescale [1]. At the same time, they can act as an interface with DC networks, enabling the operation of hybrid AC/DC systems on both the transmission and the distribution level [2, 3]. For these reasons, modern software for power system stability analysis needs to adapt to the evolving landscape introduced by IBRs.
@@ I-B Discussion on RMS vs EMT Modal Analysis
In traditional power systems, the slow dynamics of SGs were considered the most relevant, and potentially most hazardous for the proper power system operation [4]. These dynamics were the driving force behind the most common undesirable phenomena, such as local and inter-area oscillations between generators and loss of synchronism during transients [5]. The electromagnetic dynamics of the generators and the network, typically concerning frequencies much higher than the nominal one, were considered sufficiently damped and thus, not threatening to the stability of the system. For this reason, these dynamics were simplified as algebraic equations in the models used for Time Domain Simulation (TDS) and stability analysis, resulting in the classical Differential Algebraic Equation (DAE) formulation for the power system. This modeling approach is known as Root Mean Square (RMS), phasor-domain, or quasi-static phasor approach [6, 7].
The integration of IBRs into the power system has triggered new instability phenomena at various timescales, referred to as converter-driven instabilities [8]. These stability issues are commonly attributed to electromagnetic interactions between the power converter controllers, the converter filters, and the electrical grid components, such as lines, transformers, or passive filters. For this reason, detailed modeling of IBRs is required [9]. In this context, the standard quasi-stationary approximation for network and electrical elements is no longer accurate, requiring a fully Ordinary Differential Equation (ODE) representation for both network and shunt devices. This approach is termed  Electromagnetic Transient (EMT) modeling and recently has been the standard approach for TDS studies on the device level for IBR integration studies [7]. It should be noted that in the literature, different levels of detail exist between RMS and EMT modeling [6, 7].
Small-Signal Analysis (SSA) is the most frequently used method for the stability analysis of power systems during small disturbances around their nominal equilibrium point [5, 10]. The majority of commercially available software tools for power system stability analysis use RMS models for SSA, while the adoption of EMT models for stability oriented, system level SSA is still largely underdeveloped. In the scope of this work, small-signal analysis performed with RMS models will be referred to as SSA-RMS, while the one considering EMT models will be referred to as SSA-EMT.
@@ I-C Literature Review
Software solutions for power system simulation and stability analysis are typically distinguished between commercial and open-source [11]. Commercial software packages widely used in industry include PSCAD, PSS/E and Digsilent Powerfactory [12, 13, 14]. Their main scope regards TDS, for which a wide variety of models, ranging from RMS and EMT to switching models, is included in their integrated libraries. However, with regards to SSA, the included models for the transmission network revert to an RMS formulation, thus neglecting high frequency modes that may introduce adverse interactions for the power system safe operation.
Open-source software packages are usually developed for research and educational purposes. While they lack the warranties and elaborate user interface of commercial options, their libraries and algorithmic routines are typically available for inspection and modification, allowing more flexibility to individual users. Several research software projects regarding power system simulation and stability analysis have been presented in recent years [15, 16, 11, 17, 18, 19, 20]. Software packages RAMSES and ParaEMT offer solutions for efficient TDS of large power systems using parallelization techniques, based on RMS and EMT modeling of the power system, respectively [15, 16]. However, no option for SSA is included. The PSAT, implemented in Matlab environment and its follow-up Dome, implemented in Python, offer features for TDSSSA, optimal power flow and continuation power flow [11, 17]. ANDES is another Python-based library that can perform power flow calculations, TDS and SSA [18]. However, despite the option found in Dome for ad hoc inclusion of selected network dynamics [21], all three of the these software packages are based on an RMS formulation of the power system, preserving in principle the dynamic properties of the various shunt devices connected to the network, but neglecting the dynamic effect of the network itself.
Recently, software tools for the EMT-SSA have been presented [19, 20]. CSTEP, developed in Matlab environment, models the electrical element of the network as dynamic variables and additionally includes a feature for automatically removing redundant variables, for example variables representing current of inductors connected in series [19]. However, it relies on a symbolic variable internal engine which may limit the scalability of the SSA for very large systems. Reference [20] also introduced a Matlab-based, SSA toolbox, considering important effects such as transmission lines and SG stator dynamics. However neither of the two EMT-based, SSA tools offer the capability of including AC/DC hybrid systems.
@@ I-D Contributions
To the extent of the authors’ knowledge, STAMP is the first publicly available software tool for SSA-EMT of large scale hybrid AC/DC power systems, aimed for research and educational purposes [22]. Its general structure is depicted in Fig. 1. In the following, the core functionalities of STAMP are presented and its capability for SSA of converter-based systems is demonstrated through case studies based on power systems of different size and topology.
@@ II-A Power System Dynamic Modeling
A power system can be modeled as a set of ODEs [5]:
@@ II-B RMS power system modeling
with $\boldsymbol{f_{g}}$ in this formulation representing the algebraic constraints of the system and $\boldsymbol{x_{g}}$ being algebraic variables. The DAE representation, also known as RMS model, is traditionally used for the study of electromechanical and slow control dynamics up to tenths of Hz. The linear model of the system can be obtain from (3) following the process described in [23, 24].
@@ II-C EMT linear modeling
By considering $d\boldsymbol{x_{d}}/dt\neq 0$ and $d\boldsymbol{x_{g}}/dt\neq 0$ in (2), the full dynamics of the system, including the transmission grid dynamics, are included in the formulation. The linear models for SSA-EMT are obtained by linearizing the complete ODEs from (1). In particular, by applying a first-order Taylor series expansion to (1) and selecting the output variables and grouping them in the column vector $\boldsymbol{y}$, a linear state-space model of the system is obtained as follows [25]:
@@ II-D Reference Frame Transformations
Three-phase balanced power systems can be expressed in $dq$ or $qd$ rotating reference frames after applying a Park transformation, with the latter being the preferred option in STAMP [26]. Representation in rotating reference frames allows linearizing around constant operation points, simplifying the SSA. Global and local $qd$ reference frames are defined for the interconnection of several devices to the electrical grid. The global reference frame refers to a common $QD$ reference frame, defined by the selected slack device of the electrical grid, typically an ideal voltage source or a generator device. The local reference frame is the $qd-i$ reference frame from each device $i$, usually SG or Voltage Source Converter (VSC) devices.
In order to transform the vector of signals $\boldsymbol{e}_{qd-i}=[e_{q}\;e_{d}]^{\intercal}$ from the local reference frame $qd-i$ of each device $i$ to a vector $\boldsymbol{e}_{QD}=[e_{Q}\;e_{D}]^{\intercal}$ expressed in the global reference frame $QD$, rotation matrices are used as in [27, 28]:
@@ II-E Impedance representation
For Single-Input-Single-Output (SISO) systems, the connection between the state-space representation and the Transfer Function (TF) between the system input and output in the Laplace domain can be derived as [25]:
@@
The impedance and admittance TF matrices can also be expressed in the positive-negative sequence frame, which in STAMP is implemented via the transformation [29]:
@@ III Small-signal analysis methods
The STAMP implements automatized SSA considering an SSA-EMT approach representing dynamics up to a few kilohertz, but neglecting the converter switching dynamics. Despite not lying within its main scope, SSA-RMS models can also be derived from the SSA-RMS by applying various reduction techniques, for example Kron reduction [30].
@@ III-A Modal Analysis
The analysis of matrix $\boldsymbol{A}$ from (4) provides several relevant indices for the stability assessment with the most notable ones being the eigenvalues and the Participation Factors [5]. The eigenvalues $\lambda_{i}$ are calculated as the solutions of equation [10]:
@@
where ||$\cdot$|| ⋅ | is the determinant of a matrix, $i$ is the order of matrix $\boldsymbol{A}$ and $\boldsymbol{I_{i}}$ is a unity matrix of order $i$. The sign of the eigenvalues’ real part determines the stability of the system, with the dynamic system being asymptotically stable if all eigenvalues have a negative real part. In case of oscillatory modes, which are represented as complex conjugate eigenvalues, the frequency and damping ratio terms can also be obtained.
The PFs $p_{kj}$ relate the different modes of the system (represented via the eigenvalues $\lambda_{i}$) to the states of the linearised system $\Delta\boldsymbol{x}$ [23], and are calculated as [5]:
@@ III-B Impedance Analysis
The frequency-domain analysis provides information regarding the dynamic interactions between devices and the grid, and can be used to assess the overall stability of interconnected systems. Several methods have been proposed in the literature for assessing stability using impedance representations, with the most widely adopted being the impedance-based stability analysis, also known as impedance ratio analysis [31].
This method partitions the power system into two subsystems: the device under study, with admittance $\boldsymbol{Y_{d}}(s)$, and the remainder of the grid, with impedance $\boldsymbol{Z_{g}}(s)$. The resulting impedance ratio $\boldsymbol{Z_{g}}(s)\cdot\boldsymbol{Y_{d}}(s)$, also referred to as the minor loop gain, serves as the system’s open-loop TF and is used to assess stability of the system’s closed-loop. First, potential resonances are identified at the frequencies where the impedance TFs exhibits sharp peaks. Then, the stability can be evaluated by applying the Nyquist Criterion to the impedance ratio  [32]. For the cases of three-phase power systems formulated in rotating reference frames, a MIMO variation of the Nyquist Criterion, called the Generalized Nyquist Criterion (GNC), is used [32, 33]. It must be noted that to apply this approach, both subsystems must be individually stable.
@@ III-C Passivity Analysis
Assessing the passivity property of a dynamic system is another way to derive information regarding the dynamic operation of a system [34]. Specifically, the interconnection of passive elements is guaranteed to be stable, since all the dynamic elements of the network exhibit damping capability of destabilizing oscillations. The passivity of a dynamic device can be assessed across a desired frequency range $\omega_{rng}$ through its impedance TF $\boldsymbol{Z}(s)$ (equivalently through its admittance TF $\boldsymbol{Y}(s)$). This is achieved by checking the following conditions [35, 36]:
$\bullet\;\boldsymbol{Z}(s)\;\text{is stable.}$. (13)
@@ IV Implementation of state-space model
    The different subsystems are arranged and interconnected based on their respective inputs and outputs, considering the process described in [37, 38]. This subsystem interconnection creates a meshed structure of state-space subsystems that represents the linear dynamic behavior of the full system, as illustrated in Fig. 2.
The interconnection process is implemented in STAMP using the Matlab command connect from the Signal Processing Toolbox [39]. Each individual linear subsystem is grouped within higher level subsystems that serve as an intermediate level before the full model interconnection. Specifically, these groups are the following:
@@ IV-A AC grid model
$\boldsymbol{C}$ $=\begin{bmatrix}0_{2\times 1}\end{bmatrix};$ $\boldsymbol{D}$ =[±$=\begin{bmatrix}\pm 1&0&\cdots&\pm 1&0\\ 0&\pm 1&\cdots&0&\pm 1\end{bmatrix};$
where the dimension of the $\boldsymbol{D}$ matrix depends on the number of elements connected to this node, including both lines and shunt elements. The sign of the elements (±$\pm 1$) depends on the current direction of each element. In STAMP, the adopted current convention considers as positive the current injection into the node-interconnector.
@@
$m_{ij}=\begin{cases}+1&\text{if branch }j\text{ leaves bus }i,\\ -1&\text{if branch }j\text{ enters bus }i,\\ 0&\text{if branch }j\text{ is not connected to bus }i.\end{cases}$ (15)
@@ IV-B DC grid model
The DC network is modeled following the same methodology as the AC network. However, instead of using $\pi$-section line models, the DC case employs a representation consisting of three parallel $RL$ branches, which provides a more accurate representation of the DC line dynamics [40, 41, 42].
Apart from DC lines, other important elements in AC/DC systems are HVDC converters, also known as Interconnecting Power Converters [43]. These devices act as an interface between the AC and DC systems and are modeled as hybrid AC/DC admittances $Y_{\rm{ac/dc}}$, seen from both AC and DC grids. These hybrid admittances are formulated as [44]:
@@
$Y_{\rm{ac/dc}}$ can be derived from the IPC state-space model, according to the process described in Section II-E or from impedance estimation methods. The expected outputs of the IPC (inputs to the AC and DC systems) are the IPC currents ($i_{q}$, $i_{d}$, and $i_{\rm{dc}}$), and the expected inputs for the IPC element are the voltages from the AC and DC systems ($v_{q}$, $v{{}_{d}}$, and $v_{\rm{dc}}$). STAMP provides complete models of IPCs in state-space format with different control roles [45].

### v2

@@ I-A Motivation
Modern power systems are experiencing a deep transformation as a result of the massive integration of Inverter-Based Resourcess. Power electronics are being used massively in industrial and domestic loads, renewable generation units (mainly solar photovoltaics, onshore and offshore wind), Battery Energy Storage Systems (BESS), Flexible AC transmission systems (FACTS), and High-Voltage-Direct-Current (HVDC) transmission systems. Power electronics-based systems have different characteristics compared to Synchronous Generators, introducing dynamic phenomena on a much faster timescale [milano2018foundations]. At the same time, they can act as an interface with DC networks, enabling the operation of hybrid AC/DC systems on both the transmission and the distribution level [van2016hvdc, harrison2024review]. For these reasons, modern software for power system stability analysis needs to adapt to the evolving landscape introduced by IBRs.
@@ I-B Discussion on RMS vs EMT Modal Analysis
In traditional power systems, the slow dynamics of SGs were considered the most relevant, and potentially most hazardous for the proper power system operation [kundur2004definition]. These dynamics were the driving force behind the most common undesirable phenomena, such as local and inter-area oscillations between generators and loss of synchronism during transients [kundur]. The electromagnetic dynamics of the generators and the network, typically concerning frequencies much higher than the nominal one, were considered sufficiently damped and thus, not threatening to the stability of the system. For this reason, these dynamics were simplified as algebraic equations in the models used for Time Domain Simulation (TDS) and stability analysis, resulting in the classical Differential Algebraic Equation (DAE) formulation for the power system. This modeling approach is known as Root Mean Square (RMS), phasor-domain, or quasi-static phasor approach [lacerda2023phasor, lara2023revisiting].
The integration of IBRs into the power system has triggered new instability phenomena at various timescales, referred to as converter-driven instabilities [hatziargyriou2020definition]. These stability issues are commonly attributed to electromagnetic interactions between the power converter controllers, the converter filters, and the electrical grid components, such as lines, transformers, or passive filters. For this reason, detailed modeling of IBRs is required [paolone2020fundamentals]. In this context, the standard quasi-stationary approximation for network and electrical elements is no longer accurate, requiring a fully Ordinary Differential Equation (ODE) representation for both network and shunt devices. This approach is termed  Electromagnetic Transient (EMT) modeling and recently has been the standard approach for TDS studies on the device level for IBR integration studies [lara2023revisiting]. It should be noted that in the literature, different levels of detail exist between RMS and EMT modeling [lacerda2023phasor, lara2023revisiting].
Small-Signal Analysis (SSA) is the most frequently used method for the stability analysis of power systems during small disturbances around their nominal equilibrium point [kundur, milano2020eigenvalue]. The majority of commercially available software tools for power system stability analysis use RMS models for SSA, while the adoption of EMT models for stability oriented, system level SSA is still largely underdeveloped. In the scope of this work, small-signal analysis performed with RMS models will be referred to as SSA-RMS, while the one considering EMT models will be referred to as SSA-EMT.
@@ I-C Literature Review
Software solutions for power system simulation and stability analysis are typically distinguished between commercial and open-source [milano2005open]. Commercial software packages widely used in industry include PSCAD, PSS/E and Digsilent Powerfactory [PSCAD2023, PSSE2023, DIgSILENT2023]. Their main scope regards TDS, for which a wide variety of models, ranging from RMS and EMT to switching models, is included in their integrated libraries. However, with regards to SSA, the included models for the transmission network revert to an RMS formulation, thus neglecting high frequency modes that may introduce adverse interactions for the power system safe operation.
Open-source software packages are usually developed for research and educational purposes. While they lack the warranties and elaborate user interface of commercial options, their libraries and algorithmic routines are typically available for inspection and modification, allowing more flexibility to individual users. Several research software projects regarding power system simulation and stability analysis have been presented in recent years [aristidou2013dynamic, xiong2023paraemt, milano2005open, milano2013python, cui2020hybrid, serrano2022stability, kelada2025revisiting]. Software packages RAMSES and ParaEMT offer solutions for efficient TDS of large power systems using parallelization techniques, based on RMS and EMT modeling of the power system, respectively [aristidou2013dynamic, xiong2023paraemt]. However, no option for SSA is included. The PSAT, implemented in Matlab environment and its follow-up Dome, implemented in Python, offer features for TDSSSA, optimal power flow and continuation power flow [milano2005open, milano2013python]. ANDES is another Python-based library that can perform power flow calculations, TDS and SSA [cui2020hybrid]. However, despite the option found in Dome for ad hoc inclusion of selected network dynamics [milano2016semi], all three of the these software packages are based on an RMS formulation of the power system, preserving in principle the dynamic properties of the various shunt devices connected to the network, but neglecting the dynamic effect of the network itself.
Recently, software tools for the EMT-SSA have been presented [serrano2022stability, kelada2025revisiting]. CSTEP, developed in Matlab environment, models the electrical element of the network as dynamic variables and additionally includes a feature for automatically removing redundant variables, for example variables representing current of inductors connected in series [serrano2022stability]. However, it relies on a symbolic variable internal engine which may limit the scalability of the SSA for very large systems. Reference [kelada2025revisiting] also introduced a Matlab-based, SSA toolbox, considering important effects such as transmission lines and SG stator dynamics. However neither of the two EMT-based, SSA tools offer the capability of including AC/DC hybrid systems.
@@ I-D Contributions
To the extent of the authors’ knowledge, STAMP is the first publicly available software tool for SSA-EMT of large scale hybrid AC/DC power systems, aimed for research and educational purposes [STAMP_Github]. Its general structure is depicted in Fig. 1. In the following, the core functionalities of STAMP are presented and its capability for SSA of converter-based systems is demonstrated through case studies based on power systems of different size and topology.
@@ II-A Power System Dynamic Modeling
A power system can be modeled as a set of ODEs [kundur]:
@@ II-B RMS power system modeling
with $\boldsymbol{f_{g}}$ in this formulation representing the algebraic constraints of the system and $\boldsymbol{x_{g}}$ being algebraic variables. The DAE representation, also known as RMS model, is traditionally used for the study of electromechanical and slow control dynamics up to tenths of Hz. The linear model of the system can be obtain from (3) following the process described in [moutevelis2024modal, milano2010power].
@@ II-C EMT linear modeling
By considering $d\boldsymbol{x_{d}}/dt\neq 0$ and $d\boldsymbol{x_{g}}/dt\neq 0$ in (2), the full dynamics of the system, including the transmission grid dynamics, are included in the formulation. The linear models for SSA-EMT are obtained by linearizing the complete ODEs from (1). In particular, by applying a first-order Taylor series expansion to (1) and selecting the output variables and grouping them in the column vector $\boldsymbol{y}$, a linear state-space model of the system is obtained as follows [ogata_book]:
@@ II-D Reference Frame Transformations
Three-phase balanced power systems can be expressed in $dq$ or $qd$ rotating reference frames after applying a Park transformation, with the latter being the preferred option in STAMP [egea2012active]. Representation in rotating reference frames allows linearizing around constant operation points, simplifying the SSA. Global and local $qd$ reference frames are defined for the interconnection of several devices to the electrical grid. The global reference frame refers to a common $QD$ reference frame, defined by the selected slack device of the electrical grid, typically an ideal voltage source or a generator device. The local reference frame is the qdiqd-i reference frame from each device $i$, usually SG or Voltage Source Converter (VSC) devices.
In order to transform the vector of signals $\boldsymbol{e}_{qd-i}=[e_{q}\;e_{d}]^{\intercal}$ from the local reference frame qdiqd-i of each device $i$ to a vector $\boldsymbol{e}_{QD}=[e_{Q}\;e_{D}]^{\intercal}$ expressed in the global reference frame $QD$, rotation matrices are used as in [moutevelis2022bifurcation, pogaku2007modeling]:
@@ II-E Impedance representation
For Single-Input-Single-Output (SISO) systems, the connection between the state-space representation and the Transfer Function (TF) between the system input and output in the Laplace domain can be derived as [ogata_book]:
@@
The impedance and admittance TF matrices can also be expressed in the positive-negative sequence frame, which in STAMP is implemented via the transformation [rygg2016modified]:
@@ III Small-signal analysis methods
The STAMP implements automatized SSA considering an SSA-EMT approach representing dynamics up to a few kilohertz, but neglecting the converter switching dynamics. Despite not lying within its main scope, SSA-RMS models can also be derived from the SSA-RMS by applying various reduction techniques, for example Kron reduction [dorfler2012kron].
@@ III-A Modal Analysis
The analysis of matrix $\boldsymbol{A}$ from (4) provides several relevant indices for the stability assessment with the most notable ones being the eigenvalues and the Participation Factors [kundur]. The eigenvalues $\lambda_{i}$ are calculated as the solutions of equation [milano2020eigenvalue]:
@@
where $|\cdot|$ is the determinant of a matrix, $i$ is the order of matrix $\boldsymbol{A}$ and $\boldsymbol{I_{i}}$ is a unity matrix of order $i$. The sign of the eigenvalues’ real part determines the stability of the system, with the dynamic system being asymptotically stable if all eigenvalues have a negative real part. In case of oscillatory modes, which are represented as complex conjugate eigenvalues, the frequency and damping ratio terms can also be obtained.
The PFs $p_{kj}$ relate the different modes of the system (represented via the eigenvalues $\lambda_{i}$) to the states of the linearised system $\Delta\boldsymbol{x}$ [moutevelis2024modal], and are calculated as [kundur]:
@@ III-B Impedance Analysis
The frequency-domain analysis provides information regarding the dynamic interactions between devices and the grid, and can be used to assess the overall stability of interconnected systems. Several methods have been proposed in the literature for assessing stability using impedance representations, with the most widely adopted being the impedance-based stability analysis, also known as impedance ratio analysis [sun2011impedance].
This method partitions the power system into two subsystems: the device under study, with admittance $\boldsymbol{Y_{d}}(s)$, and the remainder of the grid, with impedance $\boldsymbol{Z_{g}}(s)$. The resulting impedance ratio $\boldsymbol{Z_{g}}(s)\cdot\boldsymbol{Y_{d}}(s)$, also referred to as the minor loop gain, serves as the system’s open-loop TF and is used to assess stability of the system’s closed-loop. First, potential resonances are identified at the frequencies where the impedance TFs exhibits sharp peaks. Then, the stability can be evaluated by applying the Nyquist Criterion to the impedance ratio  [amin2019nyquist]. For the cases of three-phase power systems formulated in rotating reference frames, a MIMO variation of the Nyquist Criterion, called the Generalized Nyquist Criterion (GNC), is used [amin2019nyquist, moutevelis2024virtual]. It must be noted that to apply this approach, both subsystems must be individually stable.
@@ III-C Passivity Analysis
Assessing the passivity property of a dynamic system is another way to derive information regarding the dynamic operation of a system [harnefors2015passivity]. Specifically, the interconnection of passive elements is guaranteed to be stable, since all the dynamic elements of the network exhibit damping capability of destabilizing oscillations. The passivity of a dynamic device can be assessed across a desired frequency range $\omega_{rng}$ through its impedance TF $\boldsymbol{Z}(s)$ (equivalently through its admittance TF $\boldsymbol{Y}(s)$). This is achieved by checking the following conditions [arevalo2024converter, ortega2021pid]:
$\bullet\;\boldsymbol{Z}(s)\;\text{is stable.}$ (13)
@@ IV Implementation of state-space model
    The different subsystems are arranged and interconnected based on their respective inputs and outputs, considering the process described in [nasa, gaba1988comparative]. This subsystem interconnection creates a meshed structure of state-space subsystems that represents the linear dynamic behavior of the full system, as illustrated in Fig. 2.
The interconnection process is implemented in STAMP using the Matlab command connect from the Signal Processing Toolbox [MATLAB2023b_SPT]. Each individual linear subsystem is grouped within higher level subsystems that serve as an intermediate level before the full model interconnection. Specifically, these groups are the following:
@@ IV-A AC grid model
$\boldsymbol{C}$ $=\begin{bmatrix}0_{2\times 1}\end{bmatrix}$; $\boldsymbol{D}$ $=\begin{bmatrix}\pm 1&0&\cdots&\pm 1&0\\ 0&\pm 1&\cdots&0&\pm 1\end{bmatrix}$;
where the dimension of the $\boldsymbol{D}$ matrix depends on the number of elements connected to this node, including both lines and shunt elements. The sign of the elements ($\pm 1$) depends on the current direction of each element. In STAMP, the adopted current convention considers as positive the current injection into the node-interconnector.
@@
[…]
@@ IV-B DC grid model
The DC network is modeled following the same methodology as the AC network. However, instead of using $\pi$-section line models, the DC case employs a representation consisting of three parallel $RL$ branches, which provides a more accurate representation of the DC line dynamics [Cable_DC0, Cable_DC1, Cable_DCf].
Apart from DC lines, other important elements in AC/DC systems are HVDC converters, also known as Interconnecting Power Converters [principles]. These devices act as an interface between the AC and DC systems and are modeled as hybrid AC/DC admittances $Y_{\rm{ac/dc}}$, seen from both AC and DC grids. These hybrid admittances are formulated as [hybrid_admittance]:
@@
$Y_{\rm{ac/dc}}$ can be derived from the IPC state-space model, according to the process described in Section II-E or from impedance estimation methods. The expected outputs of the IPC (inputs to the AC and DC systems) are the IPC currents ($i_{q}$, $i_{d}$, and $i_{\rm{dc}}$), and the expected inputs for the IPC element are the voltages from the AC and DC systems ($v_{q}$, $v{{}_{d}}$, and $v_{\rm{dc}}$). STAMP provides complete models of IPCs in state-space format with different control roles [arevalo2024scheduling].